# Kinetic theory of decentralized learning for smart active matter

### v1

@@ Abstract
which boil down to $\sigma^{2}(t)=[\lambda_{0}(\tau_{0}+t)]^{-1}$ and $\mu(t)=D_{\theta,T}-\mu_{0}/(\tau_{0}+t)$ for $D_{M}\to 0$. The parameter $\lambda_{0}=4\lambda_{E}\alpha_{T}\rho{V_{E}^{\prime}}^{2}$ is a characteristic rate, where $V_{E}^{\prime}={\partial\bar{V}_{x}(D_{\theta})}/{\partial D_{\theta}}|_{D_{\theta}=D_{\theta,E}}$ is the response of the average state to policy changes. The parameters $\tau_{0}$ and $\mu_{0}$ depend on initial conditions, and $D_{\theta,E}=\mathcal{P}_{E}$ is the predefined policy used in Eq. (7). When $D_{M}>0$, $\sigma^{2}(t)$ converges to a plateau value $\sigma^{2}_{\infty}=\sqrt{2D_{M}/\lambda_{0}}$ when $t\to\infty$, while $\mu(t)$ converges to the optimal policy $D_{\theta,T}$. The convergence of both $\sigma^{2}(t)$ and $\mu(t)$ occurs over the learning time $\tau_{L}=(2D_{M}\lambda_{0})^{-1/2}$. Importantly, our theoretical analysis reveals two important control parameters: the rate $\lambda_{0}$ and the mutation strength $D_{M}$. The effect of the mutation rate $D_{M}$ is characterized by the ‘uncertainty relation’ $\tau_{L}\sigma^{2}_{\infty}=\lambda_{0}^{-1}$, where both $\tau_{L}$ and $\sigma^{2}_{\infty}$ depend on $D_{M}$, while $\lambda_{0}$ is independent of $D_{M}$. The uncertainty relation states that quick convergence and adaptability (i.e., a small $\tau_{L}$) obtained by increasing the mutation rate $D_{M}$ comes at the cost of having large fluctuations around the target policy (i.e., a large $\sigma^{2}_{\infty}$), in turn implying collective fluctuations around the target state, since $\mathrm{Var}(N^{-1}\sum_{i=1}^{N}V_{x,i})\propto\sigma^{2}_{\infty}/N$. Predictions of the kinetic theory are successfully compared to agent-based simulations in Fig. 2. Equation (12) can be used to fit the long-time behavior of $\sigma^{2}(t)$ on agent-based simulations data (see Fig. 2) to determine $\lambda_{0}$ and $D_{M}$, as on experimental data where such parameters are unknown. We get $D_{M}^{\rm fit}=0.094\pm 0.002$ and $\lambda_{0}^{\rm fit}=(2.65\pm 0.01)\cdot 10^{-4}$, which compare well to the agent-based model values $D_{M}=0.1$ and $\lambda_{0}=2.655\cdot 10^{-4}$.
@@
Figure 3: Adaptation of the microswimmer model to a space-dependent target $V_{T}(x)=0.85+0.05\sin(2\pi x/L)$ (black-dashed line). Results obtained by kinetic theory (KT) and agent-based simulations (AS) are shown for different mutation rates $D_{M}.$
@@
We now consider a space-dependent target velocity $V_{T}(x)=V_{T}^{s}+V^{\Delta}_{T}\sin(2\pi x/L)$ in a system of size $L$ with periodic boundary conditions. The reward function thus also becomes space-dependent, $\mathcal{R}(\mathcal{M},x)=1-\big{(}\mathcal{M}-V_{T}(x)\big{)}^{2}$. At leading order, the steady-state average policy $\mu(x)$ satisfies [70],
$-V_{T}^{s}\nabla\mu(x)-\sqrt{2D_{M}\lambda_{L}}(\mu(x)-D_{\theta,T}(x))=0,$ (14)
@@
with a phase $\phi$ given by $\tan(\phi)=2\pi V_{T}^{s}/(\sqrt{2D_{M}\lambda_{0}}L)$ and an amplitude $V^{\Delta}_{x}(\phi)=V^{\Delta}_{T}\cos(\phi)$. Agents with small mutation rates travel across the system faster than they can adapt to the local target $V_{T}(x)$, therefore adopting an average speed $V_{T}^{s}.$ In Fig. 3, we compare this predicted behavior by kinetic theory to agent-based simulations and find very good agreement for both the mutation-rate dependent amplitude $V^{\Delta}_{x}(D_{M})$ and the phase $\phi(D_{M})$. Our results emphasize the importance of the mutation rate $D_{M}$ to maintain diversity and therefore being able to continuously adapt to space- and time-dependent targets.
@@
Figure 4: Adaptation of the light-sensing robot model to maximize collected light: Kinetic theory (KT) and agent-based simulations (AS) without mutations ($D_{M}=0$) and with mutations ($D_{M}=10^{-3}$). (a) Average policy $\mu(t)=\langle\chi\rangle$; (b) Average reward ¯$\bar{\mathcal{R}}(\chi)$ for robots with a given policy $\chi$.
@@
Light-sensing robots – We also study a model featuring robots ($d=1$) which attempt to maximize light collected from a heterogeneous light field $I(x)$. The state of a robot is $\mathcal{S}=(I,v)$, where $I$ is the local light intensity measured by a sensor, and $v$ is the robot velocity (diffusion is neglected, $D_{0}=0$). The memory $\mathcal{M}$ averages the measured light intensity according to Eq. (1), with $G(\mathcal{S})=I$. A state-dependent controller adapts the robot’s speed $v(I)=v_{0}-\chi I$, keeping $v(I)>v_{\rm min}$. The parameter $\chi$ thus defines the sensitivity with which the robot reacts to the external light. Robots adapt the policy $\mathcal{P}=\chi$ to maximize the collected light, using a reward $\mathcal{R}(\mathcal{M})=\mathcal{M}$. Figure 4(a) shows that kinetic theory successfully describes the optimization of the controller (theoretical calculations in SM [70]), by predicting the long-time behavior of $\mu(t)$. In the absence of mutations ($D_{M}=0$), the diversity $\sigma^{2}(t)$ goes to zero and $\mu(t)$ approaches the maximal reward around $\chi_{T}=1.09$ [see Fig. 4(b)]. However, for $D_{M}>0$, mutations cause the robots to learn a distribution of policies with $\sigma^{2}_{\infty}>0$. Importantly, the underlying model implies a pronounced asymmetry ¯$\bar{\mathcal{R}}(\chi_{T}+\epsilon)>\bar{\mathcal{R}}(\chi_{T}-\epsilon),\epsilon>0$ [Fig. 4(b)], where ¯$\bar{\mathcal{R}}(\chi)=\mathcal{R}(\mu_{\mathcal{M}}(\chi))$ is the average steady-state reward for an agent with policy $\chi$. This makes it preferable to adapt for $t\to\infty$ to $\mu_{\infty}>\chi_{T}$ in the presence of mutations, as observed in the agent-based simulations. Expanding $\mu_{\mathcal{M}}$ to fourth order in Eq. (7), our kinetic theory provides a quantitative prediction of $\mu_{\infty}$ [Fig. 4(a)].
@@
Conclusion and outlook – We proposed a kinetic theory framework for decentralized learning in smart active matter, in which agents maximize their reward $\mathcal{R}$ by learning improved policies $\mathcal{P}$ from neighboring agents, to reach a target collective state. This approach enables the derivation of space-dependent field equations describing the time-dependence of policy distributions. This framework has been applied to two explicit models, and predictions favorably compared to numerical simulations of the agent-based model, already showing a rich phenomenology. Future research direction may include the analysis of learning with different interaction ranges, multi-dimensional policies and more complex reward functions. While we considered for simplicity a limit of separated time scales between physical, memory and learning dynamics, our framework allows for systematic improvements over this approximation, notably taking into account the effect of the communication range. In addition, it would also be of interest to investigate the generality of the uncertainty relation linking learning time and policy fluctuations.
@@ Acknowledgements.
Acknowledgments – The authors are grateful to Olivier Dauchot for stimulating discussions. This work has been supported by MIAI@Grenoble Alpes, (ANR-19-P3IA-0003).

### v2

@@ Abstract
which boil down to $\sigma^{2}(t)=[\lambda_{0}(\tau_{0}+t)]^{-1}$ and $\mu(t)=D_{\theta,T}{+}\eta_{0}/(\tau_{0}+t)$ for $D_{\rm mut}\to 0$. The parameter $\lambda_{0}=4\lambda_{T}\alpha_{T}\rho{V_{*}^{\prime}}^{2}$ is a characteristic rate, where $V_{*}^{\prime}={\partial\bar{V}_{x}(D_{\theta})}/{\partial D_{\theta}}|_{D_{\theta}=D_{\theta,*}}$ is the response of the average velocity to policy changes and $\bar{V}_{x}(D_{\theta})=\langle V_{x}\rangle_{D_{\theta}}$ is the ensemble-averaged flow for fixed policy $D_{\theta}.$ The parameters $\tau_{0}$ and $\eta_{0}$ depend on initial conditions, and $D_{\theta,*}=\mathcal{P}_{*}$ is the predefined policy used in Eq. (7). When $D_{\rm mut}>0$, $\sigma^{2}(t)$ converges to a plateau value $\sigma^{2}_{\infty}=\sqrt{2D_{\rm mut}/\lambda_{0}}$ when $t\to\infty$, while $\mu(t)$ converges to the optimal policy $D_{\theta,T}$. The convergence of both $\sigma^{2}(t)$ and $\mu(t)$ occurs over the learning time $\tau_{L}=(2D_{\rm mut}\lambda_{0})^{-1/2}$. Importantly, our theoretical analysis reveals two important control parameters: the rate $\lambda_{0}$ and the mutation strength $D_{\rm mut}$. The effect of the mutation rate $D_{\rm mut}$ is characterized by the ‘uncertainty relation’ $\tau_{L}\sigma^{2}_{\infty}=\lambda_{0}^{-1}$, where both $\tau_{L}$ and $\sigma^{2}_{\infty}$ depend on $D_{\rm mut}$, while $\lambda_{0}$ is independent of $D_{\rm mut}$. The uncertainty relation states that quick convergence and adaptability (i.e., a small $\tau_{L}$) obtained by increasing the mutation rate $D_{\rm mut}$ comes at the cost of having large fluctuations around the target policy (i.e., a large $\sigma^{2}_{\infty}$), in turn implying fluctuations around the collective goal, since $\mathrm{Var}(N^{-1}\sum_{i=1}^{N}V_{x,i})\propto\sigma^{2}_{\infty}/N$. Predictions of the kinetic theory are successfully compared to agent-based simulations in Fig. 2. Additionally, Eq. (12) can be used to fit the long-time behavior of $\sigma^{2}(t)$ on agent-based simulations data (see Fig. 2) to determine $\lambda_{0}$ and $D_{\rm mut}$, as on experimental data where such parameters are unknown. We extract $D_{\rm mut}^{\rm fit}=0.094\pm 0.002$ and $\lambda_{0}^{\rm fit}=(2.65\pm 0.01)\cdot 10^{-4}$, which compare well to the agent-based model values $D_{\rm mut}=0.1$ and $\lambda_{0}=2.655\cdot 10^{-4}$.
@@
Figure 3: Adaptation of the microswimmer model to a space-dependent target $V_{T}(x)=0.85+0.05\sin(2\pi x/L)$ (black-dashed line). Results obtained by kinetic theory (KT) and agent-based simulations (AS) are shown for different mutation rates $D_{\rm mut}.$
@@
We now consider a space-dependent target velocity $V_{T}(x)=V_{T}^{s}+V^{\Delta}_{T}\sin(2\pi x/L)$ in a system of size $L$ with periodic boundary conditions. The reward function thus also becomes space-dependent, $\mathcal{R}(\mathcal{M},x)=-\big{(}\mathcal{M}-V_{T}(x)\big{)}^{2}$. At leading order, the steady-state average policy $\mu(x)$ satisfies [74],
$-V_{T}^{s}\nabla\mu(x)-\sqrt{2D_{\rm mut}\lambda_{0}}(\mu(x)-D_{\theta,T}(x))=0,$ (14)
@@
with a phase $\phi$ given by $\tan(\phi)=2\pi V_{T}^{s}/(\sqrt{2D_{\rm mut}\lambda_{0}}L)$ and an amplitude $V^{\Delta}_{x}(\phi)=V^{\Delta}_{T}\cos(\phi)$. The theory thus predicts that agents with smaller mutation rates have a larger phase $\phi$ and therefore the observed deviation from the target profile $\langle V_{x}(x)\rangle-V_{T}(x)=V_{T}^{\Delta}\big{(}\cos(\phi)\sin(2\pi x/L-\phi)-\sin(2\pi x/L)\big{)}$ increases. In Fig. 3, we compare this predicted behavior by kinetic theory to agent-based simulations and find very good agreement for both the mutation-rate dependent amplitude $V^{\Delta}_{x}(D_{\rm mut})$ and the phase $\phi(D_{\rm mut})$. The observed phenomenology can be explained by comparing the typical advection time $\tau_{A}=L/V_{T}^{s}$ to the learning time $\tau_{L}.$ Agents with small mutation rates have $\tau_{L}\gg\tau_{A}$ thus they learn too slow and are unable to adapt in time to the environment. Our results therefore emphasize the importance of mutations to maintain diversity and thus being able to adapt to space- or time-dependent targets.
@@
Figure 4: Adaptation of the light-sensing robot model to maximize collected light: Kinetic theory (KT) and agent-based simulations (AS) without mutations ($D_{\rm mut}=0$) and with mutations ($D_{\rm mut}=10^{-3}$). (a) Average policy $\mu(t)=\langle\chi\rangle$; (b) Average reward ¯$\bar{\mathcal{R}}(\chi)$ for robots with a given policy $\chi$.
@@
Light-sensing robots – We also study a model featuring robots (in $d=1$) which attempt to maximize light collected from a heterogeneous light field $I(x)$. The state of a robot is $\mathcal{S}=(I,v)$, where $I$ is the local light intensity measured by a sensor, and $v$ is the robot velocity (diffusion is neglected, $D_{0}=0$). The memory $\mathcal{M}$ averages the measured light intensity according to Eq. (1), with $G(\mathcal{S})=I$. A state-dependent controller adapts the robot’s speed $v(I)=v_{0}-\chi I$, keeping $v(I)\geq v_{\rm min}$. The parameter $\chi$ thus defines the sensitivity with which robots react to the external light. Robots adapt their policy $\mathcal{P}=\chi$ to maximize collected light, using a reward $\mathcal{R}(\mathcal{M})=\mathcal{M}$. Fig. 4(a) shows that the theory successfully describes the optimization of $\chi$, by predicting the long-time behavior of $\mu(t)$ (calculations in SM [74]). In the absence of mutations ($D_{\rm mut}=0$), the diversity $\sigma^{2}(t)$ goes to zero and $\mu(t)$ approaches the maximal reward around $\chi_{T}=1.09$ [see Fig. 4(b)]. However, for $D_{\rm mut}>0$, mutations cause the robots to learn a distribution of policies with $\sigma^{2}_{\infty}>0$. Importantly, the underlying model implies a pronounced asymmetry ¯$\bar{\mathcal{R}}(\chi_{T}+\epsilon)>\bar{\mathcal{R}}(\chi_{T}-\epsilon),\epsilon>0$ in the reward function [Fig. 4(b)]. Here, ¯$\bar{\mathcal{R}}(\chi)\coloneqq\mathcal{R}(\mu_{\mathcal{M}}(\chi))$ is the average steady-state reward for an agent with policy $\chi$, which we can derive exactly since the assumed time scale separation implies the existence of a direct relationship between the policy and the steady-state average memory, $\mu_{\mathcal{M}}(\chi)$. This asymmetry makes it preferable to adapt for $t\to\infty$ to $\mu_{\infty}>\chi_{T}$ in the presence of mutations, as observed in the agent-based simulations. Expanding $\mu_{\mathcal{M}}$ to fourth order in Eq. (7), KT provides a quantitative prediction of $\mu_{\infty}$ [Fig. 4(a)]. The increased steepness of ¯$\bar{\mathcal{R}}(\chi)$ for $\chi>1.2$ in KT implies a larger $V_{*}^{\prime}=\partial\bar{\mathcal{R}}(\chi)/\partial\chi$ and thus explains the faster adaptation speed observed for KT in Fig. 4(a).
@@
Conclusion and outlook – We proposed a kinetic theory framework for decentralized learning in smart active matter, in which agents maximize their reward $\mathcal{R}$ by learning improved policies $\mathcal{P}$ from neighboring agents, to reach a target collective behavior. Our theory enables derivation of space-dependent field equations describing the time-dependence of policy distributions. Assumptions required to derive closed equations include weak correlations between agents, a hierarchy of time scales, a Gaussian approximation for the $\mathcal{M}$- and $\mathcal{P}$-dependence, and expansions of the teaching probability $p_{T}$ and the average memory. Most of these assumptions can be lifted for specific applications, however, the theory cannot quantitatively account for many-body interactions and correlations. We have applied the framework to two explicit models, and predictions favorably compared to numerical simulations of the agent-based model, showing a rich phenomenology. Future research may include the analysis of learning with different interaction ranges, multi-dimensional or discrete policies and more complex reward functions such as time-dependent or non-monotonic functions or multiple species with competing reward. More generally, it will be crucial to explore systematic techniques guiding the design of reward functions facilitating targeted collective behavior. Addressing this inverse problem may be possible by incorporating higher-level learning mechanisms which are able to iterativelly optimize the agent-based reward function itself. While we considered for simplicity a limit of separated time scales between physical, memory and learning dynamics, our framework allows for systematic improvements over this approximation, notably taking into account the effect of the communication range. In addition, it would also be of interest to investigate the generality of the uncertainty relation linking learning time and policy fluctuations.
@@ Acknowledgements.
Acknowledgments – The authors are grateful to Olivier Dauchot and Bahram Houchmandzadeh for stimulating discussions. This work has been supported by a grant from MIAI@Grenoble Alpes and the Agence Nationale de la Recherche under France 2030 with the reference ANR-23-IACL-0006.